# Rank Two Sheaves With Low Discriminant on the Fano Threefold of Index 2 and Degree 5

### v1

@@
Abstract.
We describe rank 2 Gieseker semistable sheaves $E$ on the Fano threefold $X_{5}$ of index 2 and degree 5 with maximal third Chern class $c_{3}(E)$ for all possible low values of discriminant $\overline{\Delta}_{H}(E)\leq 40$. The work uses the theory of tilt-stability and Bridgeland stability conditions on smooth projective threefolds. We also make a conjecture about the rank 2 sheaves on $X_{5}$ with maximal $c_{3}$ and discriminant big enough.
The study has been funded within the framework of the HSE University Basic Research Program.
1. Introduction
The first general results on boundedness of the Chern class $c_{3}$ of rank 2 sheaves on a Fano threefold were obtained by R. Hartshorne in the case of $\mathbb{P}^{3}$ [11]. In this work he obtained exact upper bounds on $c_{3}$ of stable reflexive rank 2 sheaves on $\mathbb{P}^{3}$ [11, Theorem 8.2 (b), (d)]. In 2018 B. Schmidt proved that the bounds of Hartshorne hold true for any Gieseker semistable rank 2 sheaves without an assumption of reflexivity and described semistable sheaves with maximal $c_{3}$ [23]. His work used the theory of tilt-stability and Bridgeland stability in the bounded derived category $\mathrm{D}^{b}(\mathbb{P}^{3})$. The sheaves $E$ with maximal $c_{3}$, except for several special cases of low discriminant, were described as extensions
(1) $0\to\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\to E\to\mathcal{O}_{V}(m)\to 0$
@@ 1. Introduction
(3) $\mathrm{ch}(E)=\mathrm{rk}(E)+c_{1}H+\left(\frac{c_{1}^{2}}{2}-\frac{c_{2}}{5}\right)H^{2}+\left(\frac{c_{1}^{3}}{6}-\frac{c_{1}c_{2}}{10}+\frac{c_{3}}{10}\right)H^{3}.$
@@
In fact, in Theorem 3.4 we obtained the given bounds on the Chern classes of objects, which are $\nu_{\alpha,\beta}$-semistable for $(\beta,\alpha)$ belonging to the closure of some open set, but Proposition 2.7 implies that Gieseker semistable sheaves are tilt-semistable everywhere above the largest wall. Theorem 3.4 is formulated in terms of Chern characters, which are related to Chern classes by formula (3). Comparing the above result with the earlier works, we see that the bundle $\mathcal{U}$ from (1.1) is aCM. The sheaves from (2.3), when they are locally free, are at the same time aCM, instanton in sense of [12], and twisted by $\mathcal{O}_{X}(-1)$ Ulrich bundles.
@@
2. The Derived Category $\mathrm{D}^{b}(X)$ and Stability
@@
Recall that the slope of a coherent sheaf $E\in\operatorname{Coh}(X)$ is defined as $\mu(E)=\frac{H^{2}\cdot\mathrm{ch}_{1}(E)}{H^{3}\cdot\mathrm{ch}_{0}(E)}$ (in case of division by 0 we set $\mu(E)=+\infty$). A coherent sheaf $E$ is called $\mu$-stable (resp. $\mu$-semistable) or slope stable (resp. slope semistable) if for any proper subsheaf $0\neq F\hookrightarrow E$ the inequality $\mu(F)<\mu(E/F)$ (resp. $\mu(F)\leq\mu(E/F)$) holds.
Following [4] and [23], we obtain a new heart of a bounded t-structure on $\mathrm{D}^{b}(X)$ using the process of tilting. In general, let $\mathcal{A}$ be the heart of a bounded t-structure on a triangulated category $\mathfrak{D}$. Recall that a torsion pair is a pair $(\mathcal{T},\mathcal{F})$ of subcategories of the category $\mathcal{A}$ such that, firstly, $\operatorname{Hom}(T,F)=0$ for any $T\in\mathcal{T},F\in\mathcal{F}$ and, secondly, for any $E\in\mathcal{A}$ there is a short exact sequence $0\to T\to E\to F\to 0$ in $\mathcal{A}$ with $T\in\mathcal{T},F\in\mathcal{F}$.
@@
$\mathcal{T}_{\beta}=\{E\in\operatorname{Coh}(X)\colon\mathrm{any}\;\mathrm{quotient}\;E\to G\;\mathrm{satisfies}\;\mu(G)>\beta\},$
$\mathcal{F}_{\beta}=\{E\in\operatorname{Coh}(X)\colon\mathrm{any}\;\mathrm{subsheaf}\;F\to E\;\mathrm{satisfies}\;\mu(F)\leq\beta\}.$
In general, if $(\mathcal{T},\mathcal{F})$ is a torsion pair on the heart $\mathcal{A}$ as defined above, its tilt is the extension closure $\langle\mathcal{F}[1],\mathcal{T}\rangle\subset\mathfrak{D}$. It is also the heart of a bounded t-structure. In our case the new heart $\operatorname{Coh}^{\beta}(X)$ is the extension closure $\langle\mathcal{F}_{\beta}[1],\mathcal{T}_{\beta}\rangle$. Equivalently, $\operatorname{Coh}^{\beta}(X)$ can be described as the full additive subcategory of the derived category, consisting of those complexes $E\in\mathrm{D}^{b}(X)$ for which $\mathcal{H}^{-1}(E)\in\mathcal{F}_{\beta},\mathcal{H}^{0}(E)\in\mathcal{T}_{\beta}$ and $\mathcal{H}^{i}(E)=0$ for $i\neq-1,0$.
@@
Let $\alpha>0$ be a positive real number. Following Schmidt ([22, 24]), for an object $E\in\operatorname{Coh}^{\beta}(X)$ we define a central charge function
@@
which is a $\mathbb{Z}$-linear map $Z^{\mathrm{tilt}}_{\alpha,\beta}:K(\mathrm{D}^{b}(X))=K(\operatorname{Coh}^{\beta}(X))\to\mathbb{C}$.
@@
Proposition 2.1.
For any $\nu_{\alpha,\beta}$-semistable object $E\in\operatorname{Coh}^{\beta}(X)$ we have
(4) $\overline{\Delta}_{H}(E):=(H^{3})^{2}\cdot Q^{\mathrm{tilt}}(E)=(H^{2}\cdot\mathrm{ch}_{1}(E))^{2}-2(H^{3}\cdot\mathrm{ch}_{0}(E))(H\cdot\mathrm{ch}_{2}(E))=\\ =(H^{2}\cdot\mathrm{ch}_{1}^{\beta}(E))^{2}-2(H^{3}\cdot\mathrm{ch}_{0}^{\beta}(E))(H\cdot\mathrm{ch}_{2}^{\beta}(E))\geq 0.$
@@
$\mathcal{T}^{\prime}_{\alpha,\beta}=\{E\in\operatorname{Coh}^{\beta}(X)\colon\text{any quotient}\,E\twoheadrightarrow G\;\mathrm{satisfies}\;\nu_{\alpha,\beta}(G)>0\},$
$\mathcal{F}^{\prime}_{\alpha,\beta}=\{E\in\operatorname{Coh}^{\beta}(X)\colon\text{any non-trivial subobject}\,F\hookrightarrow E\;\mathrm{satisfies}\;\nu_{\alpha,\beta}(F)\leq 0\},$
@@
It is known that for $s>\frac{1}{6}$ $(\mathcal{A}^{\alpha,\beta}(X),Z_{\alpha,\beta,s})$ is a Bridgeland stability condition [14, Corollary 0.2]. This result follows from the fact that the following generalized Bogomolov-Gieseker inequality holds for $X$ ([14, Theorem 0.1], [3, Theorem 5.4]):
Proposition 2.2.
@@
Fix $v\in\Lambda=\mathbb{Z}^{2}\oplus\frac{1}{10}\mathbb{Z}$. There is a locally finite wall-and-chamber structure in the upper half-plane $\mathbb{H}:=\{(\beta,\alpha)\in\mathbb{R}^{2}\ |\ \alpha>0\}$ such that for an object $E\in\mathrm{Coh}^{\beta}(X)$ with $\mathrm{ch}_{\leq 2}(E)=v$ tilt-stability of $E$ does not change when $(\beta,\alpha)$ vary within a chamber ([3, Proposition B.5]). A numerical wall with respect to $v\in\Lambda$ is a non-trivial proper subset $W$ of $\mathbb{H}$ given by an equation of the form $\nu_{\alpha,\beta}(v)=\nu_{\alpha,\beta}(w)$ for another element $w\in\Lambda$. We denote this numerical wall by $W(v,w)$. A numerical wall $W$ is called an actual wall (or simply a wall) if the set of semistable objects with class $v$ changes at $W$. The structure of walls in tilt-stability is well understood. We have the following Proposition.
Proposition 2.3 (Numerical properties of walls).
@@
  5. (5)
[…]
Proof.
The statements (1)-(3) were proven by Schmidt [24, Theorem 3.3] for the case of $X=\mathbb{P}^{3}$. In that case $v,w\in\mathbb{Z}^{2}\oplus\frac{1}{2}\mathbb{Z}$, but the calculations can be done for any $v,w\in\mathbb{R}^{3}$. The statement (4) follows from [24, Lemma 7.3]. The statement (5) for the case of $\mathbb{P}^{3}$ is [23, Lemma 2.10], and the calculation can be done in our case too (we only need to multiply the vector $(v_{1},2v_{2},3v_{3})$ by 5 so that $5(v_{1},2v_{2},3v_{3})\in\Lambda$). ∎
@@
([3, Proposition B.5]). Following Schmidt, we denote a wall given by such an equation by $W(E,F)$. If $W(E,F)$ is semicircular, we denote its radius by $\rho(E,F)$. We will also use the properties of walls described in the following Proposition.
Proposition 2.4 (Properties of actual walls).
@@
Proof.
The statement (1) is proven in [17, Lemma 2.4]. The statement (2) is essentially proven in [24, Lemma 2.7]: it suffices to note that in the case of a semicircular wall we have $\mathrm{ch}_{\leq 2}(F)\neq 0,\mathrm{ch}_{\leq 2}(G)\neq 0$ and $\mathrm{ch}_{\leq 2}(F)-\lambda\mathrm{ch}_{\leq 2}(G)\neq 0$ for any $\lambda\in\mathbb{R}$. Hence, if we apply this lemma with the quadratic form $Q(E)=\overline{\Delta}_{H}(E)$ defined on $\Lambda\otimes\mathbb{R}$, we get that for some $\lambda>0$ and $(\beta,\alpha)\in W$ $Z^{\mathrm{tilt}}_{\alpha,\beta}(\mathrm{ch}_{\leq 2}(F)-\lambda\mathrm{ch}_{\leq 2}(G))=0$, so $Q(\mathrm{ch}_{\leq 2}(F)-\lambda\mathrm{ch}_{\leq 2}(G))<0$, hence $Q(F,G):=Q^{\mathrm{tilt}}(\mathrm{ch}_{\leq 2}(F),\mathrm{ch}_{\leq 2}(G))>0$ (we use the Bogomolov-Gieseker inequality for $F$ and $G$) and $Q(E)=Q(F)+Q(G)+2Q(F,G)>Q(F)+Q(G)$.
@@
Finally, the statement (5) appeared in the proof of [5, Lemma 3.6] and was formulated in [25, Proposition 2.7 (vi)]. ∎
@@
Proof.
This is proven in [4, Lemma 7.2.1, Lemma 7.2.2]. Note that, if we assume that $H^{2}\cdot\mathrm{ch}_{1}^{\beta}(E)$ has the minimal positive value and that $E$ is $\nu_{\alpha,\beta}$-semistable, then it is $\nu_{\alpha,\beta}$-stable. ∎
Remark 2.6.
Proposition 2.4 (2) implies that objects $E\in\operatorname{Coh}^{\beta}(X)$ with $\overline{\Delta}_{H}(E)=0$ can only be destabilized at the vertical wall. Hence, by Lemma 2.5 line bundles and their shifts by one are tilt-semistable everywhere (under the condition that they belong to $\operatorname{Coh}^{\beta}(X)$).
We will also use the following relation between tilt-stability and the more classical notions of stability. Let $f,g\in\mathbb{R}[m]$ be polynomials. If $\deg(f)<\deg(g)$, then we set $f>g$. If $\deg(f)=\deg(g)$ and $a$, $b$ are the leading coefficients in $f$, $g$ respectively, then we put $f<(\leq)\ g$ if $\frac{f(m)}{a}<(\leq)\ \frac{g(m)}{b}$ for all $m\gg 0$. For an arbitrary sheaf $E\in\operatorname{Coh}(X)$ we define the numbers $a_{i}(E)$ for $i\in\{0,1,2,3\}$ through the Hilbert polynomial $P(E,m):=\chi(E(m))=a_{3}(E)m^{3}+a_{2}(E)m^{2}+a_{1}(E)m+a_{0}(E)$. In addition, we set $P_{2}(E,m):=a_{3}(E)m^{2}+a_{2}(E)m+a_{1}(E)$. The sheaf $E\in\operatorname{Coh}(X)$ is called Gieseker (semi)stable, if for any its proper subsheaf $0\neq F\hookrightarrow E$ the inequality $P(F,m)<(\leq)\ P(E/F,m)$ holds. Respectively, $E$ is called (semi)stable if for any subsheaf $0\neq F\hookrightarrow E$ the inequality $P_{2}(F,m)<(\leq)\ P_{2}(E/F,m)$ holds. Gieseker stability, 2-stability and $\mu$-stability of a sheaf satisfy the relations:
@@
Proof.
This is proven in [6, Prop. 14.2] for K3 surfaces, but the proof works in our case too. ∎
@@
Proof.
This is a special case of [5, Theorem 3.1] (applied to E(nH)E(-nH)) with $H^{3}=5,i_{X}=2$.∎
@@
Lemma 2.9.
Let $E\in\mathrm{Coh}^{\beta}(X)$ be a tilt-stable object with $\mathrm{ch}_{0}(E)>0$. If there is a semicircular wall $W$ for $E$, then $W$ is induced by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{0}(F)>0,\mu(F)<\mu(E).$
Proof.
Proposition 2.4 implies that $W$ is induced by a subobject or quotient $F^{\prime}$ with $\mathrm{ch}_{0}(F^{\prime})>0,\mu(F^{\prime})<\mu(E);$ the objects $E$ and $F^{\prime}$ should be tilt-semistable at $W$. Take $(\beta,\alpha)\in W$. If $F^{\prime}$ is not $\nu_{\alpha,\beta}$-stable, we can consider its Jordan–Hölder filtration in $\nu_{\alpha,\beta}$-stability. So, if $F^{\prime}$ is a subobject of $E$, then we take $F$ to be the subobject of $F^{\prime}$, which is $\nu_{\alpha,\beta}$-stable with $\nu_{\alpha,\beta}(F)=\nu_{\alpha,\beta}(F^{\prime})$; respectively, if $F^{\prime}$ is a quotient of $E$, we take $F$ to be the quotient of $F^{\prime}$, which is $\nu_{\alpha,\beta}$-stable with $\nu_{\alpha,\beta}(F)=\nu_{\alpha,\beta}(F^{\prime})$. Note that $F$ also defines the wall $W$. The object $E$ is tilt-stable on one side of $W$, therefore both $F^{\prime}$ and $F$ destabilize $E$ on the other side. It means that $\frac{\mathrm{ch}_{0}(F^{\prime})}{\mathrm{ch}_{1}^{\beta}(F^{\prime})}-\frac{\mathrm{ch}_{0}(E)}{\mathrm{ch}_{1}^{\beta}(E)}$ and $\frac{\mathrm{ch}_{0}(F)}{\mathrm{ch}_{1}^{\beta}(F)}-\frac{\mathrm{ch}_{0}(E)}{\mathrm{ch}_{1}^{\beta}(E)}$ have the same sign. Together with the fact that $\beta<\mu(E)$, it implies that $\mu(F)<\mu(E),\mathrm{ch}_{0}(F)>0$. ∎

### v2

@@
UDC 512.7
Rank Two Sheaves With Low Discriminant on the Fano Threefold of Index 2 and Degree 5
D. A. Vassiliev [^1] The study has been funded within the framework of the HSE University Basic Research Program.
Аннотация
In this paper rank 2 Gieseker semistable sheaves $E$ on the Fano threefold $X_{5}$ of index 2 and degree 5 with maximal third Chern class $c_{3}(E)$ for all possible low values of discriminant $\overline{\Delta}_{H}(E)\leq 40$ are described. The work uses the theory of tilt-stability and Bridgeland stability conditions on smooth projective threefolds. Also a conjecture about the rank 2 sheaves on $X_{5}$ with maximal $c_{3}$ and discriminant big enough is proposed.
Keywords: Fano threefolds, semistable coherent sheaves, Bridgeland stability conditions.
1 Introduction
The first general results on boundedness of the Chern class $c_{3}$ of rank 2 sheaves on a Fano threefold were obtained by R. Hartshorne in the case of $\mathbb{P}^{3}$ [10]. In this work he obtained exact upper bounds on $c_{3}$ of stable reflexive rank 2 sheaves on $\mathbb{P}^{3}$ [10, Theorem 8.2 (b), (d)]. In 2018 B. Schmidt proved that the bounds of Hartshorne hold true for any Gieseker semistable rank 2 sheaves without an assumption of reflexivity and described semistable sheaves with maximal $c_{3}$ [23]. His work used the theory of tilt-stability and Bridgeland stability in the bounded derived category $\mathrm{D}^{b}(\mathbb{P}^{3})$. The sheaves $E$ with maximal $c_{3}$, except for several special cases of low discriminant, were described as extensions
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\to E\to\mathcal{O}_{V}(m)\to 0$ (1)
@@ 1 Introduction
$\mathrm{ch}(E)=\mathrm{rk}(E)+c_{1}H+\left(\frac{c_{1}^{2}}{2}-\frac{c_{2}}{5}\right)H^{2}+\left(\frac{c_{1}^{3}}{6}-\frac{c_{1}c_{2}}{10}+\frac{c_{3}}{10}\right)H^{3}.$ (3)
@@
In fact, in Theorem 3.4 we obtained the given bounds on the Chern classes of tilt-semistable objects, but Proposition 2.7 implies that Gieseker semistable sheaves are tilt-semistable everywhere above the largest wall. Theorem 3.4 is formulated in terms of Chern characters, which are related to Chern classes by formula (3). Comparing the above result with the earlier works, we see that the bundle $\mathcal{U}$ from (1.1) is aCM. The sheaves from (2.3), when they are locally free, are at the same time aCM, instanton in sense of [12], and twisted by $\mathcal{O}_{X}(-1)$ Ulrich bundles.
@@
2 The Derived Category $\mathrm{D}^{b}(X)$ and Stability
@@
Recall that the slope of a coherent sheaf $E\in\mathrm{Coh}(X)$ is defined as $\mu(E)=\frac{H^{2}\cdot\mathrm{ch}_{1}(E)}{H^{3}\cdot\mathrm{ch}_{0}(E)}$ (in case of division by 0 we set $\mu(E)=+\infty$). A coherent sheaf $E$ is called $\mu$-stable (resp. $\mu$-semistable) or slope stable (resp. slope semistable) if for any proper subsheaf $0\neq F\hookrightarrow E$ the inequality $\mu(F)<\mu(E/F)$ (resp. $\mu(F)\leq\mu(E/F)$) holds.
Following [3] and [23], we obtain a new heart of a bounded t-structure on $\mathrm{D}^{b}(X)$ using the process of tilting. In general, let $\mathcal{A}$ be the heart of a bounded t-structure on a triangulated category $\mathfrak{D}$. Recall that a torsion pair is a pair $(\mathcal{T},\mathcal{F})$ of subcategories of the category $\mathcal{A}$ such that, firstly, $\mathrm{Hom}(T,F)=0$ for any $T\in\mathcal{T},F\in\mathcal{F}$ and, secondly, for any $E\in\mathcal{A}$ there is a short exact sequence $0\to T\to E\to F\to 0$ in $\mathcal{A}$ with $T\in\mathcal{T},F\in\mathcal{F}$.
@@
$\mathcal{T}_{\beta}=\{E\in\mathrm{Coh}(X)\colon\mathrm{any}\;\mathrm{quotient}\;E\to G\;\mathrm{satisfies}\;\mu(G)>\beta\},$
$\mathcal{F}_{\beta}=\{E\in\mathrm{Coh}(X)\colon\mathrm{any}\;\mathrm{subsheaf}\;F\to E\;\mathrm{satisfies}\;\mu(F)\leq\beta\}.$
In general, if $(\mathcal{T},\mathcal{F})$ is a torsion pair on the heart $\mathcal{A}$ as defined above, its tilt is the extension closure $\langle\mathcal{F}[1],\mathcal{T}\rangle\subset\mathfrak{D}$. It is also the heart of a bounded t-structure. In our case the new heart $\mathrm{Coh}^{\beta}(X)$ is the extension closure $\langle\mathcal{F}_{\beta}[1],\mathcal{T}_{\beta}\rangle$. Equivalently, $\mathrm{Coh}^{\beta}(X)$ can be described as the full additive subcategory of the derived category, consisting of those complexes $E\in\mathrm{D}^{b}(X)$ for which $\mathcal{H}^{-1}(E)\in\mathcal{F}_{\beta},\mathcal{H}^{0}(E)\in\mathcal{T}_{\beta}$ and $\mathcal{H}^{i}(E)=0$ for $i\neq-1,0$.
@@
Let $\alpha>0$ be a positive real number. Following Schmidt ([22, 24]), for an object $E\in\mathrm{Coh}^{\beta}(X)$ we define a central charge function
@@
which is a $\mathbb{Z}$-linear map $Z^{\mathrm{tilt}}_{\alpha,\beta}:K(\mathrm{D}^{b}(X))=K(\mathrm{Coh}^{\beta}(X))\to\mathbb{C}$.
@@
Proposition 2.1
For any $\nu_{\alpha,\beta}$-semistable object $E\in\mathrm{Coh}^{\beta}(X)$ we have
$\overline{\Delta}_{H}(E):=(H^{3})^{2}\cdot Q^{\mathrm{tilt}}(E)=(H^{2}\cdot\mathrm{ch}_{1}(E))^{2}-2(H^{3}\cdot\mathrm{ch}_{0}(E))(H\cdot\mathrm{ch}_{2}(E))=\\ =(H^{2}\cdot\mathrm{ch}_{1}^{\beta}(E))^{2}-2(H^{3}\cdot\mathrm{ch}_{0}^{\beta}(E))(H\cdot\mathrm{ch}_{2}^{\beta}(E))\geq 0.$ (4)
@@
$\mathcal{T}^{\prime}_{\alpha,\beta}=\{E\in\mathrm{Coh}^{\beta}(X)\colon\text{any quotient}\,E\twoheadrightarrow G\;\mathrm{satisfies}\;\nu_{\alpha,\beta}(G)>0\},$
$\mathcal{F}^{\prime}_{\alpha,\beta}=\{E\in\mathrm{Coh}^{\beta}(X)\colon\text{any non-trivial subobject}\,F\hookrightarrow E\;\mathrm{satisfies}\;\nu_{\alpha,\beta}(F)\leq 0\},$
@@
It is known that for $s>\frac{1}{6}$ $(\mathcal{A}^{\alpha,\beta}(X),Z_{\alpha,\beta,s})$ is a Bridgeland stability condition [14, Corollary 0.2]. This result follows from the fact that the following generalized Bogomolov-Gieseker inequality holds for $X$ ([14, Theorem 0.1], [2, Theorem 5.4]):
Proposition 2.2
@@
Fix $v\in\Lambda=\mathbb{Z}^{2}\oplus\frac{1}{10}\mathbb{Z}$. There is a locally finite wall-and-chamber structure in the upper half-plane $\mathbb{H}:=\{(\beta,\alpha)\in\mathbb{R}^{2}\ |\ \alpha>0\}$ such that for an object $E\in\mathrm{Coh}^{\beta}(X)$ with $\mathrm{ch}_{\leq 2}(E)=v$ tilt-stability of $E$ does not change when $(\beta,\alpha)$ vary within a chamber ([2, Proposition B.5]). A numerical wall with respect to $v\in\Lambda$ is a non-trivial proper subset $W$ of $\mathbb{H}$ given by an equation of the form $\nu_{\alpha,\beta}(v)=\nu_{\alpha,\beta}(w)$ for another element $w\in\Lambda$. We denote this numerical wall by $W(v,w)$. A numerical wall $W$ is called an actual wall (or simply a wall) if the set of semistable objects with class $v$ changes at $W$. The structure of walls in tilt-stability is well understood. We have the following Proposition.
Proposition 2.3 (Numerical properties of walls)
@@
  5. 5.
    If $\overline{\Delta}_{H}(v)>0$, then the set of $(\beta,\alpha)$ for which $Q_{\alpha,\beta}(v_{0},v_{1}H,v_{2}H^{2},v_{3}H^{3})=0$ is the semicircular numerical wall $W((v_{0},v_{1},v_{2}),(5v_{1},10v_{2},15v_{3}))$. The locus where $Q_{\alpha,\beta}(v_{0},v_{1}H,v_{2}H^{2},v_{3}H^{3})<0$ is the semidisk bounded by $Q_{\alpha,\beta}(v_{0},v_{1}H,v_{2}H^{2},v_{3}H^{3})=0$ and $\beta$-axis.
Proof. The statements (1)-(3) were proven by Schmidt [24, Theorem 3.3] for the case of $X=\mathbb{P}^{3}$. In that case $v,w\in\mathbb{Z}^{2}\oplus\frac{1}{2}\mathbb{Z}$, but the calculations can be done for any $v,w\in\mathbb{R}^{3}$. The statement (4) follows from [24, Lemma 7.3]. The statement (5) for the case of $\mathbb{P}^{3}$ is [23, Lemma 2.10], and the calculation can be done in our case too (we only need to multiply the vector $(v_{1},2v_{2},3v_{3})$ by 5 so that $5(v_{1},2v_{2},3v_{3})\in\Lambda$). $\square$
@@
([2, Proposition B.5]). Following Schmidt, we denote a wall given by such an equation by $W(E,F)$. If $W(E,F)$ is semicircular, we denote its radius by $\rho(E,F)$. We will also use the properties of walls described in the following Proposition.
Proposition 2.4 (Properties of actual walls)
@@
Proof. The statement (1) is proven in [17, Lemma 2.4]. The statement (2) is essentially proven in [24, Lemma 2.7]: it suffices to note that in the case of a semicircular wall we have $\mathrm{ch}_{\leq 2}(F)\neq 0,\mathrm{ch}_{\leq 2}(G)\neq 0$ and $\mathrm{ch}_{\leq 2}(F)-\lambda\mathrm{ch}_{\leq 2}(G)\neq 0$ for any $\lambda\in\mathbb{R}$. Hence, if we apply this lemma with the quadratic form $Q(E)=\overline{\Delta}_{H}(E)$ defined on $\Lambda\otimes\mathbb{R}$, we get that for some $\lambda>0$ and $(\beta,\alpha)\in W$ $Z^{\mathrm{tilt}}_{\alpha,\beta}(\mathrm{ch}_{\leq 2}(F)-\lambda\mathrm{ch}_{\leq 2}(G))=0$, so $Q(\mathrm{ch}_{\leq 2}(F)-\lambda\mathrm{ch}_{\leq 2}(G))<0$, hence $Q(F,G):=Q^{\mathrm{tilt}}(\mathrm{ch}_{\leq 2}(F),\mathrm{ch}_{\leq 2}(G))>0$ (we use the Bogomolov-Gieseker inequality for $F$ and $G$) and $Q(E)=Q(F)+Q(G)+2Q(F,G)>Q(F)+Q(G)$.
@@
Finally, the statement (5) appeared in the proof of [4, Lemma 3.6] and was formulated in [25, Proposition 2.7 (vi)]. $\square$
@@
Proof. This is proven in [3, Lemma 7.2.1, Lemma 7.2.2]. Note that, if we assume that $H^{2}\cdot\mathrm{ch}_{1}^{\beta}(E)$ has the minimal positive value and that $E$ is $\nu_{\alpha,\beta}$-semistable, then it is $\nu_{\alpha,\beta}$-stable. $\square$
Remark 2.6
Proposition 2.4 (2) implies that objects $E\in\mathrm{Coh}^{\beta}(X)$ with $\overline{\Delta}_{H}(E)=0$ can only be destabilized at the vertical wall. Hence, by Lemma 2.5 line bundles and their shifts by one are tilt-semistable everywhere (under the condition that they belong to $\mathrm{Coh}^{\beta}(X)$).
We will also use the following relation between tilt-stability and the more classical notions of stability. Let $f,g\in\mathbb{R}[m]$ be polynomials. If $\deg(f)<\deg(g)$, then we set $f>g$. If $\deg(f)=\deg(g)$ and $a$, $b$ are the leading coefficients in $f$, $g$ respectively, then we put $f<(\leq)\ g$ if $\frac{f(m)}{a}<(\leq)\ \frac{g(m)}{b}$ for all $m\gg 0$. For an arbitrary sheaf $E\in\mathrm{Coh}(X)$ we define the numbers $a_{i}(E)$ for $i\in\{0,1,2,3\}$ through the Hilbert polynomial $P(E,m):=\chi(E(m))=a_{3}(E)m^{3}+a_{2}(E)m^{2}+a_{1}(E)m+a_{0}(E)$. In addition, we set $P_{2}(E,m):=a_{3}(E)m^{2}+a_{2}(E)m+a_{1}(E)$. The sheaf $E\in\mathrm{Coh}(X)$ is called Gieseker (semi)stable, if for any its proper subsheaf $0\neq F\hookrightarrow E$ the inequality $P(F,m)<(\leq)\ P(E/F,m)$ holds. Respectively, $E$ is called (semi)stable if for any subsheaf $0\neq F\hookrightarrow E$ the inequality $P_{2}(F,m)<(\leq)\ P_{2}(E/F,m)$ holds. Gieseker stability, 2-stability and $\mu$-stability of a sheaf satisfy the relations:
@@
Proof. This is proven in [5, Prop. 14.2] for K3 surfaces, but the proof works in our case too. $\square$
@@
Proof. This is a special case of [4, Theorem 3.1] (applied to E(nH)E(-nH)) with $H^{3}=5,i_{X}=2$. $\square$
@@
Lemma 2.9
Let $E$ be a tilt-semistable object with $\mathrm{ch}_{0}(E)>0$. If there is a semicircular wall $W$ for $E$, then $W$ is induced by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{0}(F)>0,\mu(F)<\mu(E).$
Proof. Choose a point $(\beta_{0},\alpha_{0})\in W$. The wall $W$ is given by a short exact sequence
$0\to F^{\prime}\to E\to G^{\prime}\to 0$ (5)
in $\mathrm{Coh}^{\beta_{0}}(X)$. Suppose that (5) destabilizes $E$ below $W$. Choose $\alpha_{1}<\alpha_{0}$ so that the point $(\beta_{0},\alpha_{1})$ lies in the chamber adjacent to $W$. Since $E$ is not $\nu_{\alpha_{1},\beta_{0}}$-semistable, we can take a tilt-stable subobject $F\hookrightarrow E$ such that $\nu_{\alpha_{1},\beta_{0}}(F)>\nu_{\alpha_{1},\beta_{0}}(E)$. We should have $\nu_{\alpha_{0},\beta_{0}}(F)=\nu_{\alpha_{0},\beta_{0}}(E)$ and $\nu_{\alpha,\beta_{0}}(F)>\nu_{\alpha,\beta_{0}}(E)$ for $\alpha<\alpha_{0}$, hence $\frac{\mathrm{ch}_{0}(F)}{\mathrm{ch}_{1}^{\beta_{0}}(F)}>\frac{\mathrm{ch}_{0}(E)}{\mathrm{ch}_{1}^{\beta_{0}}(E)}>0$. Therefore $\frac{\mathrm{ch}_{1}^{\beta_{0}}(F)}{\mathrm{ch}_{0}(F)}<\frac{\mathrm{ch}_{1}^{\beta_{0}}(E)}{\mathrm{ch}_{0}(E)}$ and we obtain $\mu(F)<\mu(E),\mathrm{ch}_{0}(F)>0$. Finally, if (5) destabilizes $E$ above $W$, then we analogously choose a point $(\beta_{0},\alpha_{1})$ in the chamber located immediately above $W$ and take a tilt-stable quotient $E\twoheadrightarrow F$ such that $\nu_{\alpha_{1},\beta_{0}}(E)>\nu_{\alpha_{1},\beta_{0}}(F)$. $\square$